# Rope to Nope and Back Again: A New Hybrid Attention Strategy

### v1

@@ Abstract
Long context large language models (LLMs) have achieved remarkable advancements, driven by techniques like Rotary Position Embedding (RoPE) (Su et al., 2023) and its extensions (Chen et al., 2023; Liu et al., 2024c; Peng et al., 2023). By adjusting RoPE parameters and incorporating training data with extended contexts, we can train performant models with considerably longer input sequences. However, existing RoPE-based methods exhibit performance limitations when applied to extended context lengths. This paper presents a comprehensive analysis of various attention mechanisms, including RoPE, No Positional Embedding (NoPE), and Query-Key Normalization (QK-Norm), identifying their strengths and shortcomings in long context modeling. Our investigation identifies distinctive attention patterns in these methods and highlights their impact on long context performance, providing valuable insights for architectural design. Building on these findings, we propose a novel architecture based on a hybrid attention mechanism that not only surpasses conventional RoPE-based transformer models in long context tasks but also achieves competitive performance on benchmarks requiring shorter context lengths.
Large Language Model, Long-Context, Pretraining, ICML
@@ 1 Introduction
Developing language models capable of handling long context lengths poses several challenges. First, as the context length increases, an effective modeling of extended input sequences becomes increasingly critical. This often requires advancements in positional encoding (Su et al., 2023), extrapolation techniques (Ding et al., 2024), or architectural innovations (Tworkowski et al., 2023; Huang et al., 2024). Second, training long context large language models with billions of parameters demands significant computational resources. Overcoming this challenge requires scalable algorithms, high-quality datasets, and robust infrastructure. Lastly, deploying these models in real-world applications demands low latency and low memory usage, which requires meticulous optimization of both the model architecture and the serving infrastructure. Addressing these issues requires a strategic balance between modeling, training, and deployment efficiency, while managing potential trade-offs.
@@
On the modeling front, two components of the transformer architecture are particularly crucial for long context capabilities: the attention mechanism and positional embeddings. Recent research has proposed various methods to enhance these components. For instance, Landmark Attention (Mohtashami & Jaggi, 2023) trains attention modules to select relevant blocks using a representative token, referred to as a “landmark token”, for efficient retrieval within extended text corpora. Similarly, Focused Transformer (Tworkowski et al., 2023) adopts a contrastive training approach to prioritize attending the most relevant portions of the input sequence, allowing the model to focus on smaller, contextually significant subsets of tokens. Although these approaches improve long context modeling, stabilizing training remains a key challenge for extending transformer capabilities to longer sequences. Query-Key Normalization (QK-Norm) (Henry et al., 2020; Rybakov et al., 2024) has been introduced to address the stability issue, which normalizes the query-key vectors along the head dimension before computing attention. Although QK-Norm mitigates numerical instability during training and is widely used (Team, 2024; Dehghani et al., 2023; Li et al., 2024b), it may impair long context capabilities.
@@
In addition to the chosen attention mechanism, positional embeddings play a crucial role in long context modeling. Various approaches have been proposed to improve their effectiveness. Popular methods include Absolute Position Embedding (APE) (Vaswani et al., 2017), Relative Position Embedding (Raffel et al., 2023), ALiBi (Press et al., 2022), and Rotary Position Embedding (RoPE) (Su et al., 2023). Among these, RoPE has gained significant adoption in large language models (LLMs) (Dubey et al., 2024; Yang et al., 2024a; Üstün et al., 2024) due to its simplicity and effectiveness. In particular, it has the ability to extrapolate context lengths by adjusting RoPE $\theta$ values during training (Liu et al., 2024a; AI et al., 2024; Dang et al., 2024). Other techniques, such as relative bias (Press et al., 2022; Raffel et al., 2023; Chi et al., 2022) and contextualized position embeddings (Golovneva et al., 2024), introduce distance-based bias terms or condition the position information on input semantics. These methods often affect attention distribution by incorporating auxiliary information, such a positional indices or explicit recency bias. However, whether certain information or biases are beneficial to long context modeling or overall performance remains less explored. Additionally, the concept of No Positional Embedding (NoPE) has been explored by (Kazemnejad et al., 2023), suggesting that removing explicit positional embeddings and relying solely on implicit positional information derived from the causal mask can enhance long context performance.
Though the aforementioned strategies help with long context modeling, training and serving long context models still present notable challenges. The vanilla attention mechanism that most transformers use have quadratic complexity relative to sequence length, making them prohibitively expensive for extended contexts. Techniques such as Sliding Window Attention (SWA) (Jiang et al., 2023; Team et al., 2024a) limit each token’s attention to a fixed-size window of neighboring tokens, reducing computational overhead while preserving local contextual understanding. Other methods, like sparse attention (Child et al., 2019; Beltagy et al., 2020; Tay et al., 2020a), generalize SWA by applying various sparsity patterns, including random sparsity (Zaheer et al., 2021) and dilated attention (Ding et al., 2023). More recent strategies, such as activation beacon (Zhang et al., 2024), further optimize attention by compressing local keys and values sequentially, or attention sink (Xiao et al., 2024), which points out that tokens at the beginning of the sequences contribute to a disproportionately large mass of attention score and demonstrating that retaining these tokens and employing a sliding window approach can stabilize loss over longer contexts. On the serving side, methods like KV cache trimming (Zhang et al., 2023; Li et al., 2024a; Cai et al., 2024) selectively reduce KV cache size using certain heuristics, thereby minimizing inference memory usage and improving throughput. While reducing attention span can improve efficiency, it often involves trade-offs in model performance, underscoring the importance of a balanced design.
@@
In this paper, we begin analyzing attention patterns of different attention mechanisms,, NoPE, and QK-Norm and its impacts on long context performance trained up to 750 billion tokens. Building on these insights, we propose a new modeling approach and extensively pretrain up to 5 trillion tokens, followed by supervised fine-tuning on a diverse set of datasets tailored for long context. We show that this approach surpasses existing state-of-the-art extrapolation-based RoPE models (Liu et al., 2024c) by a large margin, striking a balance between efficiency and performance.
@@ 2 Observation
In this section, we assess three models with different attention components, use needles-in-a-haystack (Kamradt, 2023) (NIAH) and analyze the attention patterns to understand how these variants affect performance. This analysis guides our architectural design choices throughout this work.
@@ 2.1 Experimental Setup
All model variants share a common configuration consisting of 8 billion parameters (including the token embedding parameters), with detailed architectural specifications provided in Table 1. For this set of experiments, the model is trained in two stages: a pretraining stage followed by a supervised fine-tuning (SFT) stage. Previous research shows that the SFT stage is necessary for long context evaluations, as it can reduce variance in long context tasks and enables the emergence of long context capabilities that may not manifest in models trained solely through pretraining (Gao et al., 2024).
@@
Parameters 8B
Embedding Dim 4096
FFN Dim 28672
Non-linearity swiglu
Num Layers 32
Num Heads 32
Num KV Heads 8
Vocab Size 256000
@@
    RoPE Model: For this variant, we employ Rotary Position Embedding (RoPE) to encode positional information. During the pretraining stage, the RoPE parameter $\theta$ is set to 10,000. In the subsequent supervised fine-tuning (SFT) stage, $\theta$ is increased to 2 million to account for the increased context length. This variant serves as the baseline model configuration, maintaining an architecture similar to that of most existing models (Dubey et al., 2024; Jiang et al., 2023; Cohere For AI, 2024).
@@
    QK-Norm Model: Layer normalization (Ba et al., 2016) is applied to both the query and the key vectors before performing the angular rotation used in RoPE. All other hyperparameters, including the $\theta$ values and training methodology, remain identical to those of the RoPE variant.
@@
    NoPE Model: Previous research (Wang et al., 2024; Haviv et al., 2022) has demonstrated that transformer variants trained without positional embeddings (NoPE) can perform effectively on long context tasks. However, these models often exhibit inferior performance in terms of perplexity and downstream task evaluations within the trained sequence length (Haviv et al., 2022). In our study, the NoPE variant does not have QK-Norm. This variant is trained using the same methodology as the other two variants.
@@ 2.2.1 Evaluations
We evaluate the variants on a set of core evaluation benchmarks, including MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), CommonsenseQA (Talmor et al., 2019), ARC (Clark et al., 2018) for core model capabilities and NIAH benchmark (Kamradt, 2023) for long context capability. NIAH evaluates a model’s ability to retrieve information accurately from a specific sentence (the “needle”) embedded within a lengthy document (the “haystack”). The needle is randomly placed at varying depths within the sequence to examine performance across different context lengths. To improve robustness, we modify the original NIAH benchmark, where each context-depth combination is tested 16 times with different random seeds, creating diverse context compositions for comprehensive evaluation. The results of all standard benchmark evaluations and results with 65536 context length needles are presented in Table 2. Although prior research has emphasized the limitations of NIAH (Xu et al., 2024) for evaluating deeper and more general context understanding, our focus is solely on testing basic long context capabilities and gaining insights on model architecture design, for which this benchmark is sufficient.
Table 2 shows that the RoPE and QK-Norm variants exhibit comparable performance on standard benchmarks, while the NoPE variant lags behind. This finding is consistent with previous studies (Kazemnejad et al., 2023; Wang et al., 2024). For long context evaluations, QK-Norm performs the worst among the three variants, despite its decent performance in other capabilities. This observation is consistent with the results from the comparisons between Command R and Command R+, where Command R, despite being a significantly smaller model, outperforms Command R+ overall on longer context benchmarks (Hsieh et al., 2024). Although the NoPE variant has slightly lower needles score compared to the RoPE variant, it is decent given that its base capabilities is relatively low.
Model Val Loss MMLU HellaSwag CommonsenseQA ARC-E ARC-C Needles 65k
RoPE 1.52 48.55 73.74 68.30 81.05 39.13 9.82
QK-Norm 1.53 48.21 73.68 68.23 80.54 38.98 7.93
NoPE 1.58 47.61 72.16 66.42 76.94 37.12 9.03
Table 2: Comparison of models on a variety of benchmarks. All evaluations are based on the performance of the SFT-models.
Context Length Model Variants begin needle context qc
8k RoPE 0.3863 0.0328 0.3809 0.2000
QK-Norm 0.0242 0.0173 0.8020 0.1565
NoPE 0.3058 0.0454 0.4501 0.1987
32k RoPE 0.3541 0.0201 0.4343 0.1915
QK-Norm 0.0064 0.0056 0.8517 0.1364
NoPE 0.2807 0.0325 0.4981 0.1886
128k RoPE 0.3463 0.0010 0.4751 0.1776
QK-Norm 0.0010 0.0004 0.8993 0.0994
NoPE 0.0846 0.0073 0.8156 0.0925
Table 3: Needles Attention Pattern: RoPE, QK-Norm and Nope
@@ 2.2.2 Attention Pattern Analysis
To investigate the impacts of different architectures, we also analyze the attention patterns within each model. This approach is inspired by previous studies (Ye et al., 2024; Leviathan et al., 2024), which examined attention scores to gain insight on improving model performance.
We still utilize the NIAH task by dividing the context into four segments: the first 10 tokens (begin), the needle sentence tokens (needle), general context tokens (context), and question/completion tokens (qc). We position the needles at approximately 50% depth within the context to increase the complexity of the task, as most models suffer from the lost-in-the-middle problem, as highlighted in previous works (Liu et al., 2024b; Baker et al., 2024). For each model, we first calculate the attention scores between the query tokens of “qc” and the key tokens of all four segments across all heads and layers. The attention scores are summed within each segment and then aggregated across all heads and layers to obtain the average attention score for each segment. These scores are further averaged across multiple samples at sequence lengths of 8,000 tokens, 32,000 tokens, and 128,000 tokens. We refer to this metric as “attention mass” in the following sections. The results are reported in Table 8.
Table 8 reveals that all variants exhibit decreasing attention mass on needle tokens as the sequence length increases. This indicates that retrieving relevant information becomes more challenging as the sequence grows longer. Additionally, within each context length, the NoPE variant demonstrates the highest attention mass on needle tokens, followed by the RoPE variant, while the QK-Norm variant has the lowest attention mass. We also notice that the QK-Norm variant has extremely low attention mass for the beginning tokens (attention sink (Xiao et al., 2024)) compared to other variants and higher attention mass over the noisy context. This is consistent with QK-Norm’s relatively poor performance in the NIAH task. We argue that QK-Norm has this effect because the normalization operation mitigates magnitude information from the dot product of Query and Key vectors which tends to result in attention logits being closer in terms of magnitude and flatter in terms of distribution. A more detailed analysis can be found in Appendix B.
@@ 2.2.3 Hybrid Model
Building on the findings above, we combine RoPE and NoPE to leverage the strengths of both approaches. NoPE’s effective attention mechanism for retrieving information based on vector similarity and RoPE’s explicit modeling of positional information and recency bias can, in combination, enhance general performance. We propose a new variant that alternates between NoPE and Rotary Position Embedding (RoPE) layers. Specifically, the two position-embedding strategies are interleaved, with NoPE applied in one layer and RoPE in the next. To ensure consistency and facilitate comparisons, the RoPE parameter $\theta$ is initially set to 10,000 during pre-training. We then conduct multiple fine-tuning runs with varying $\theta$ values, including 10,000, 100,000, 2 million, and 4 million, to evaluate the model’s performance across different configurations. We refer to this variant as the RNoPE variant and RNoPE-10k, RNoPE-100k, RNoPE-2M, and RNoPE-4M with the numbers indicating the RoPE $\theta$ value used during SFT.
Model NoPE Layers RoPE Layers needles-128k
begin needle context qc begin needle context qc
RoPE - - - - 0.3541 0.0201 0.4343 0.1915 7.395
RNoPE-10k 0.3275 0.0765 0.5672 0.0287 0.0049 0.0004 0.6805 0.3142 8.036
RNoPE-100k 0.3263 0.0778 0.5633 0.0327 0.0241 0.0005 0.6782 0.2972 7.461
RNoPE-2M 0.3250 0.0712 0.5735 0.0303 0.1111 0.0046 0.6233 0.2611 7.022
RNoPE-4M 0.3486 0.0369 0.5981 0.0165 0.0960 0.0039 0.6774 0.2227 6.203
RNoPE-10k-swa 0.3303 0.0742 0.5634 0.0321 - - - - 9.562
Table 4: Needles Attention Pattern: RoPE and RNoPE variants
@@
We report the needles evaluation score at a sequence length of 128,000 and the attention mass calculations (as in Table 8) for all variants. The results are shown in Table 4. The attention mass is aggregated separately for all RoPE and NoPE layers. For simplicity, we present results based on 32,000 sequence length samples, with the complete table for all sequence lengths provided in Appendix A.
@@
Results in Table 4 reveal that as we increase the RoPE parameter $\theta$ during fine-tuning, the model’s long context capability diminishes. This finding contradicts previous observations in pure RoPE models (Men et al., 2024; Liu et al., 2024a), where larger RoPE $\theta$ value during pretraining or fine-tuning improves long context performance and expands the effective attention receptive field. To explore this discrepancy, we compare the attention mass between different model variants.
First, we observe a clear difference in behavior among the RNoPE variants layers. The NoPE layers exhibit strong retrieval capabilities, characterized by a pronounced spike in attention mass on the needle tokens and a phenomenon of attention sink (Xiao et al., 2024) at the beginning tokens. Furthermore, the NoPE layers show significantly weaker recency bias compared to pure RoPE or pure NoPE models. In contrast, the RoPE layers in RNoPE variants demonstrate extremely weak retrieval capabilities and demonstrate almost no attention sink, as evidenced by their relatively low attention mass on the beginning and needle tokens. However, these RoPE layers exhibit a much stronger recency bias compared to the pure RoPE variant or their NoPE layer counterparts.
@@
When comparing between different $\theta$ values, we find that as $\theta$ increases, the recency bias of the RoPE layers decreases, as reflected in the decreasing attention mass of the qc tokens. This supports previous findings and theoretical analyses suggesting that increasing $\theta$ expands the effective receptive field of the attention mechanism (Men et al., 2024), resulting in a flatter attention distribution. However, our empirical observations suggest that a larger receptive field in the RoPE layers introduces noise, which disrupts the ability of the subsequent NoPE layers to compute similarities and perform retrieval tasks effectively, ultimately leading to lower needle evaluation scores. For example, the needle attention mass of the NoPE layers dropped from 0.0765 to 0.0369, and the needle evaluation score decreased from 8.036 to 6.203 as $\theta$ increased from 10,000 to 4 million.
@@
    Combining NoPE and RoPE layers provides distinct advantages in processing information. NoPE layers excel at information retrieval, while RoPE layers handle local information effectively due to their built-in recency bias.
@@
    Restricting the effective attention span of the RoPE layers in RNoPE models can denoise the attention mass and ensure that each layer type adheres to its specialized role.
Based on these insights, we fine-tune a new variant, RNoPE-10k-swa, where “swa” stands for sliding window attention. This approach introduces a hard limit on the attention span of the RoPE layers, enforcing the second insight from above. Specifically, we set the sliding window size for the RoPE layers to 8,192, while maintaining the full attention span of the NoPE layers for long context retrieval. The rest of the training process remains identical to the RNoPE-10k variant, with no changes to $\theta$. The evaluation results, reported in Table 4, demonstrate a notable improvement. The RNoPE-10k-swa variant achieves a needles-128k score of 9.562, outperforming both the baseline and the RNoPE-10k model. Additionally, its NoPE layers exhibit a well-structured attention pattern, indicative of strong long context retrieval capabilities.

### v2

@@ Abstract
Long context large language models (LLMs) have achieved remarkable advancements, driven by techniques like Rotary Position Embedding (RoPE) [61] and its extensions [12, 47, 54]. By adjusting RoPE parameters and incorporating training data with extended contexts, we can train performant models with considerably longer input sequences. However, existing RoPE-based methods exhibit performance limitations when applied to extended context lengths. This paper presents a comprehensive analysis of various attention mechanisms, including RoPE, No Positional Embedding (NoPE), and Query-Key Normalization (QK-Norm), identifying their strengths and shortcomings in long context modeling. Our investigation identifies distinctive attention patterns in these methods and highlights their impact on long context performance, providing valuable insights for architectural design. Building on these findings, we propose a novel architecture featuring a hybrid attention mechanism that integrates global and local attention spans. This design not only surpasses conventional RoPE-based transformer models with full attention in both long and short context tasks but also delivers substantial efficiency gains during training and inference.
@@ 1 Introduction
Developing language models capable of handling long context lengths poses several challenges. First, as the context length increases, an effective modeling of extended input sequences becomes increasingly critical. This often requires advancements in positional encoding [61], extrapolation techniques [23], or architectural innovations [69, 34]. Second, training long context large language models with billions of parameters demands significant computational resources. Overcoming this challenge requires scalable algorithms, high-quality datasets, and robust infrastructure. Lastly, deploying these models in real-world applications demands low latency and low memory usage, which requires meticulous optimization of both the model architecture and the serving infrastructure.
@@
On the modeling front, two components of the transformer architecture are particularly crucial for long context capabilities: the attention mechanism and positional embeddings. Recent research has proposed various methods to enhance these components. For instance, Landmark Attention [53] trains attention modules to select relevant blocks using a representative token, referred to as a “landmark token", for efficient retrieval within extended text corpora. Similarly, Focused Transformer [69] adopts a contrastive training approach to prioritize attending the most relevant portions of the input sequence, allowing the model to focus on smaller, contextually significant subsets of tokens. Although these approaches improve long context modeling, stabilizing training remains a key challenge for extending transformer capabilities to longer sequences. Query-Key Normalization (QK-Norm) [31, 57] has been introduced to address the stability issue, which normalizes the query-key vectors along the head dimension before computing attention. Although QK-Norm mitigates numerical instability during training and is widely used [66, 21, 42], it may impair long context capabilities.
@@
In addition to the chosen attention mechanism, positional embeddings play a crucial role in long context modeling. Various approaches have been proposed to improve their effectiveness. Popular methods include Absolute Position Embedding (APE) [70], Relative Position Embedding [56], ALiBi [55], and Rotary Position Embedding (RoPE) [61]. Among these, RoPE has gained significant adoption in large language models (LLMs) [24, 75, 83] due to its simplicity and effectiveness. In particular, it has the ability to extrapolate context lengths by adjusting RoPE $\theta$ values during training [45, 2, 18]. Other techniques, such as relative bias [55, 56, 13] and contextualized position embeddings [26], introduce distance-based bias terms or condition the position information on input semantics. These methods often affect attention distribution by incorporating auxiliary information, such a positional indices or explicit recency bias. However, whether certain information or biases are beneficial to long context modeling or overall performance remains less explored. Additionally, the concept of No Positional Embedding (NoPE) has been explored by [38], suggesting that removing explicit positional embeddings and relying solely on implicit positional information derived from the causal mask can enhance long context performance.
Despite advances in long context modeling, training and serving such models remain challenging due to the quadratic complexity of standard attention. Techniques like Sliding Window Attention (SWA) [36, 67] mitigate this by restricting each token’s attention to a local window, reducing compute while maintaining local coherence. Sparse attention methods [14, 7, 64] extend this by introducing structured sparsity, including random [78] and dilated patterns [22]. More recent approaches further compress attention, such as activation beacon [80], which sequentially summarizes local keys and values, and attention sink [73], which stabilizes long-sequence training by preserving early tokens with a sliding window. On the serving side, KV cache trimming methods [81, 41, 10] selectively discard cached states based on heuristics to reduce memory usage and boost inference efficiency. However, these gains often come with trade-offs in model quality, emphasizing the need for careful design.
@@
In this paper, we begin analyzing attention patterns of different attention mechanisms, RoPE, NoPE, and QK-Norm and its impacts on long context performance trained up to 750 billion tokens. Building on these insights, we propose a novel hybrid attention architecture and extensively pretrain up to 5 trillion tokens, followed by supervised fine-tuning on a diverse set of datasets tailored for long context. We show that this architecture surpasses existing state-of-the-art extrapolation-based RoPE models [47] by a large margin, striking a balance between efficiency and performance.
@@ 2 Observation
In this section, we assess three models with different attention components on needles-in-a-haystack [37] (NIAH) and analyze the attention patterns to understand how these variants affect performance. Analysis in this section guides our architectural design choices throughout this work.
@@ 2.1 Experimental Setup
All model variants share a common configuration consisting of 8 billion parameters (including the token embedding parameters), with detailed architectural specifications provided in Table 1. For this set of experiments, the model is trained in two stages: a pretraining stage followed by a supervised fine-tuning (SFT) stage. Previous research shows that the SFT stage is necessary for long context evaluations, as it can reduce variance in long context tasks and enables the emergence of long context capabilities that may not manifest in models trained solely through pretraining [25].
@@
Parameters Emb. Dim FFN Dim Non-linearity Layers Heads KV Heads Vocab Size Values 4096 28672 swiglu 32 32 8 256000
@@
    RoPE Model: For this variant, we employ Rotary Position Embedding (RoPE) to encode positional information. During the pretraining stage, the RoPE parameter $\theta$ is set to 10,000. In the subsequent supervised fine-tuning (SFT) stage, $\theta$ is increased to 2 million to account for the increased context length. This variant serves as the baseline model configuration, maintaining an architecture similar to that of most existing models [24, 36, 17].
@@
    QK-Norm Model: Layer normalization [4] is applied to both the query and the key vectors before performing the angular rotation used in RoPE. All other hyperparameters, including the $\theta$ values and training methodology, remain identical to those of the RoPE variant.
@@
    NoPE Model: Previous research [72, 29] has demonstrated that transformer variants trained without positional embeddings (NoPE) can perform effectively on long context tasks. However, these models often exhibit inferior performance in terms of perplexity and downstream task evaluations within the trained sequence length [29]. In our study, the NoPE variant does not have QK-Norm. This variant is trained using the same methodology as the other two variants.
@@ 2.2.1 Evaluations
We evaluate the variants on a set of core evaluation benchmarks, including MMLU [30], HellaSwag [79], CommonsenseQA [63], ARC [15] for core model capabilities and NIAH benchmark [37] for long context capability. NIAH evaluates a model’s ability to retrieve information accurately from a specific sentence (the “needle”) embedded within a lengthy document (the “haystack”). The needle is randomly placed at varying depths within the sequence to examine performance across different context lengths. To improve robustness, we modify the original NIAH benchmark, where each context-depth combination is tested 16 times with different random seeds, creating diverse context compositions for comprehensive evaluation. The results of all standard benchmark evaluations and results with 65536 context length needles are presented in Table 2. Although prior research has emphasized the limitations of NIAH [74] for evaluating deeper and more general context understanding, our focus is solely on testing basic long context capabilities and gaining insights on model architecture design, for which this benchmark is sufficient.
Table 2 shows that the RoPE and QK-Norm variants exhibit comparable performance on standard benchmarks, while the NoPE variant lags behind. This finding is consistent with previous studies [38, 72]. For long context evaluations, QK-Norm performs the worst among the three variants, despite its decent performance in other capabilities. This observation is consistent with the results from the comparisons between Command R and Command R+, where Command R, despite being a significantly smaller model, outperforms Command R+ overall on longer context benchmarks [32]. Although the NoPE variant has slightly lower needles score compared to the RoPE variant, it is decent given that its base capabilities is relatively low.
@@ 2.2.2 Attention Pattern Analysis
To better understand the impacts of different architectures, we also analyze the attention patterns within each model. This approach is inspired by previous studies [77, 39] where attention scores assigned to different parts of the context are closely examined.
We still utilize the NIAH task by first dividing the context into four segments:
  • Begin: The first 10 tokens. This part of the context is also often referred to as the “attention sink” [73], where a disproportionately large amount of attention is typically allocated in a trained transformer model.
  • Needle: The tokens of the inserted needle sentence. Ideally, a well-trained model should assign a relatively high amount of attention to this part of the context.
  • End: The query and completion tokens, which can represent recency bias.
  • Context: The remainder of the context, typically consisting of noise or irrelevant content.
We position the needles at approximately 50% depth within the context to increase the complexity of the task, as most models suffer from the lost-in-the-middle problem, as highlighted in previous works [46, 5]. For each model, we first calculate the attention scores between the query tokens of “End” segment and the key tokens of all four segments across all heads and layers. The attention scores are summed within each segment and then aggregated across all heads and layers to obtain the average attention weight for each segment. These scores are further averaged across multiple samples at sequence lengths of 8,000 tokens, 32,000 tokens, and 128,000 tokens. We refer to this metric as “attention mass” in the following sections. The results of each variant are visualized in Figure 1.
We begin by comparing attention distributions across different sequence lengths from Figure 1. We observe a consistent decrease in attention mass allocated to the “Needle” segment and a corresponding increase in attention mass on the “Context” segment as sequence length increases. This trend indicates that retrieving relevant information becomes increasingly difficult as the context grows longer. When comparing across model variants, the NoPE variant consistently allocates the highest attention mass to the Needle” segment, followed by the RoPE variant, while the QK-Norm variant assigns the least attention to this segment. Furthermore, the QK-Norm variant exhibits markedly lower attention mass on the “Begin” segment and substantially higher attention mass on the “Context” segment relative to other variants. This indicates that models trained with the QK-Norm component exhibit a weak attention sink and are more susceptible to interference from noisy content. These patterns are consistent with QK-Norm’s poor performance on the NIAH task. We argue that QK-Norm has this effect because the normalization operation mitigates magnitude information from the dot product of Query and Key vectors which tends to result in attention logits being closer in terms of magnitude and flatter in terms of distribution. A more detailed analysis of why QK-Norm is detrimental to long-context modeling is provided in Appendix B.
Model Val Loss MMLU HellaSwag CommonsenseQA ARC-E ARC-C Needles 65k RoPE 1.52 48.55 73.74 68.30 81.05 39.13 9.82 QK-Norm 1.53 48.21 73.68 68.23 80.54 38.98 7.93 NoPE 1.58 47.61 72.16 66.42 76.94 37.12 9.03
Table 2: Comparison of model performance across a range of benchmarks. All evaluations are based on the outputs of the SFT models. Red cells indicate lower performance.
Figure 1: Attention Distribution across context lengths of each variant
@@ 2.2.3 Hybrid Model
Building on the findings above, we are inspired to combine RoPE and NoPE layer-wise to leverage the strengths of both variants. NoPE provides an effective attention mechanism for retrieving information based on vector similarity, while RoPE explicitly models positional information and introduces a recency bias. By combining them, we aim to enhance overall performance. We propose a new variant that alternates between NoPE and Rotary Position Embedding (RoPE) layers. Specifically, the two position-embedding strategies are interleaved, with NoPE applied in one layer and RoPE in the next. To ensure consistency and enable meaningful comparisons, the RoPE parameter $\theta$ is initially set to 10,000 during pre-training. The pre-training procedure follows the setup described in Section 2.1 in terms of data and training protocol. Subsequently, we perform multiple fine-tuning runs with varying $\theta$ values—10,000, 100,000, 2 million, and 4 million—to evaluate performance across different configurations, using the same training steps and data as in Section 2.1. We refer to these models as RNoPE-10k, RNoPE-100k, RNoPE-2M, and RNoPE-4M with the specific RoPE $\theta$ value used for each during supervised fine-tuning (SFT).
Figure 2: Attention Distribution at 32k context length for rope baseline and hybrid variants.
Model RoPE RNoPE-10k RNoPE-100k RNoPE-2M RNoPE-4M RNoPE-10k-swa Needles-128k Score 7.40 8.04 7.46 7.02 6.20 9.56
Table 3: Needles score at 128k for all model variants
@@
For the RoPE baseline model and all RNoPE variants, we report the needles evaluation score of all SFT models at a sequence length of 128,000 in Table 3. We also display the attention mass of all variants in Figure 2. The attention mass is aggregated separately for all RoPE and NoPE layers. For simplicity, we present results based on samples with 32,000 sequence lengths, with the complete table for all sequence lengths provided in Appendix A.
@@
The results in Table 3 reveal that increasing the RoPE parameter $\theta$ during fine-tuning of the RNoPE variants—where NoPE and RoPE layers are interleaved—leads to a decline in the model’s long-context capability. In contrast, previous studies on pure RoPE architectures [50, 45] have shown that using a larger RoPE $\theta$ during pre-training or fine-tuning enhances long-context performance and expands the effective attention receptive field. To investigate this discrepancy, we compare the attention mass across different model variants from Figure 2.
First, when comparing the NoPE and RoPE layers across all RNoPE variants, we observe a clear divergence in their behavior. The NoPE layers exhibit strong retrieval capabilities, characterized by a pronounced spike in attention mass on the “Needle” segment and a phenomenon of attention sink [73] on the “Begin” segment. Additionally, these layers show a significantly weaker recency bias indicated by the low attention mass on the “End” segment. In contrast, the RoPE layers within the RNoPE variants demonstrate extremely weak retrieval performance, evidenced by near-zero attention mass on the “Needle” segment and only modest attention on the “Begin” segment—indicating an absence of an attention sink. However, these RoPE layers exhibit a much stronger recency bias compared to the pure RoPE baseline. In summary, interleaving RoPE and NoPE layers induces a spontaneous “division of labor” phenomenon, where RoPE layers focus on local information aggregation and NoPE layers specialize in long-range retrieval. Remarkably, this functional specialization emerged naturally during training, without any explicit training constraints, data augmentation strategies, or specific loss function designs.
@@
Next, we examine the RNoPE variants fine-tuned with different $\theta$ values. As $\theta$ increases, we observe a consistent decline in the recency bias of the RoPE layers. Specifically, the attention mass on the “End” segment drops from 0.314 in RNoPE-10k to 0.223 in RNoPE-4M, indicating a flatter attention distribution that reaches deeper into the context. This aligns with prior findings and theoretical analyses suggesting that larger $\theta$ values expand the effective receptive field of the attention mechanism [50]. However, our empirical results indicate that this expanded receptive field in the RoPE layers introduces noise into the overall architecture, which disrupts the NoPE layers’ ability to compute similarities and perform retrieval effectively. This degradation is reflected in both attention mass and task performance: the needle attention mass in the NoPE layers drops from 0.0765 to 0.0369, and the needle evaluation score decreases from 8.036 to 6.203 as $\theta$ increases from 10,000 to 4 million. These findings further underscore that the model spontaneously develops a “division of labor” mechanism during training, with distinct roles emerging between RoPE and NoPE layers.
@@
    Division of Labour: Combining NoPE and RoPE layers yields complementary strengths, with each type naturally assuming specialized roles after training. NoPE layers are adapted to information retrieval, while RoPE layers become effective at modeling local context due to their inherent recency bias.
@@
    Potential Efficiency Gains: Restricting the effective attention span of the RoPE layers in RNoPE models can reduce noise in the attention distribution and reinforce the functional specialization of each layer. Additionally, it lowers the computational cost (FLOPs) during training—particularly for longer context lengths—thereby improving training efficiency while maintaining or even enhancing performance.
Guided by these insights, we fine-tune a new variant, RNoPE-10k-swa, where “swa” denotes sliding window attention. This modification imposes a hard limit on the attention span of RoPE layers, operationalizing the second insight above. Specifically, the sliding window size for RoPE layers is set to 8,192, while the NoPE layers retain full attention to support long-context retrieval. All other training configurations remain identical to the RNoPE-10k variant, including the use of $\theta=10{,}000$. Evaluation results, presented in Table 3, show a marked improvement. The RNoPE-10k-swa variant achieves a needles-128k score of 9.562, surpassing both the baseline and the original RNoPE-10k model. Moreover, its NoPE layers exhibit a well-structured attention pattern with high attention mass on both the “Begin” and “End” segments from 2, reflecting strong long-context retrieval capabilities.